Find the intervals on which the function f defind by f(x)=x^4-8x^2+16 is concave upward or concave downward.
1
Expert's answer
2020-05-15T17:29:23-0400

Given

"f(x)=x^4-8x^2+16"

Since

"f'(x) = 4x^3-16x\\
f''(x)= 12x^2-16"

Solve

"f''(x)=0"

we get

"x=\frac{2\sqrt{3}}{3},\ \ \ \ \:x=-\frac{2\sqrt{3}}{3}"

Now , we check the sign of "f''(x)" , since

"f''(-2)>0,\ \ \ f''(0)<0,\ \ f''(2)>0"

Then "f(x)" is concave up on "(-\infty , -2\sqrt{3}/3)\cup (2\sqrt{3}/3,\infty)" and

concave down on "(-2\sqrt{3}/3,2\sqrt{3}/3)"